# Seeing through Light and Darkness: Sensor-Physics Grounded Deblurring HDR NeRF from Single-Exposure Images and Events

### v1

@@ 2.2 ERGB-based novel view synthesis
Event cameras are bio-inspired sensors detecting brightness changes at individual pixels with high temporal resolution [9]. This scene radiance measurement paradigm has higher dynamic range and temporal resolution compared to traditional frame-based RGB cameras, attracting a lot of research in the fields of event-based image deblurring [29, 18, 22, 36, 43], HDR image reconstruction [38, 12, 40, 28, 41, 11, 6], and high-speed HDR video generation [33, 52, 44, 50]. Some recent works explore event-based NVS [17, 35, 24, 24, 25, 48, 10]. While other works integrate both image and event (ERGB) data into the deblurring NVS task, such as NeRF-based works [31, 30, 4, 32] and 3DGS-based works [8, 15, 46] Addressing the limitations of blurry LDR inputs, EvHDR-NeRF [5] first jointly exploits single-exposure images and events to reconstruct high-fidelity HDR radiance fields. However, these methods do not fully consider the gap between sensor output and physical radiance, leading to limited HDR and deblurring NeRF novel view synthesis performance.
@@ 5.1 Experiment settings
Synthetic dataset: The synthetic dataset consists of 8 HDR “Blender” scenes from HDR-NeRF [16]. For each scene, there are 18 training views and 17 test views. We use Blender [1] and v2e [14] to generate the blurry images at exposure time $\Delta t_{2}$ and corresponding events as in [31] for training views. We additionally provide sharp LDR images at exposure times $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ for the image-based HDR NVS methods without deblurring effects. For test views, we provide GT sharp HDR images and LDR images at $\Delta t_{1},\Delta t_{3}$ for HDR and novel exposure evaluation as in [16].
Real dataset: The real dataset includes 5 scenes captured under extreme lighting conditions using the DAVIS 346 event camera [37]. Each scene has 16 training views and 12 test views. For training views, we use a handheld DAVIS 346 to capture the blurry LDR images at exposure time $\Delta t_{2}$ along with the corresponding event data. We also captured sharp LDR images with a tripod at exposure times $\Delta t_{0},\Delta t_{2},\Delta t_{4}$. For the test views, we use a tripod to capture GT sharp LDR images at exposure times $\Delta t_{1}$, $\Delta t_{3}$. We also capture the GT sharp LDR images at $\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$, and use the classic multi-exposure HDR reconstruction algorithm DeBevec [7] to generate the GT sharp HDR image with the 5 sharp LDR images at different exposures for each test view as in RawHDR [51].
@@ 5.3 Deblurring NVS experiment
See-NeRFEvD Simple Linear [4] ours 22.02 .8669 .2474 16.43 .6932 .3210 29.52 .9267 .2776
@@ 5.4 Ablation study
Effect of event mapping field (EMF): Comparing See-NeRFNoEMF and See-NeRF in Tab. 3, our EMF marginally helps align the CRF curve (Fig. 8), achieving further improvement on HDR and Deblurring NVS. However, replacing our EMF with the simple linear model in EvDeblur-NeRF [4] leads to barely satisfactory reconstruction results and CRF estimation (See-NeRFEvD in Tab. 3 and Fig. 8).
@@ 5.5 Discussions
Robustness to LDR images with different exposures and event noise: As shown in Tab. 4, we conduct a robustness experiment on the synthetic “Catroom” scene for the HDR NVS task. Firstly, we use the single-exposure images at $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ as inputs, respectively. The results show that single-exposure inputs with longer exposure times yield better results, indicating that the low-light regions of the scene are more difficult to learn than the high-light regions. Secondly, we use events with different noise levels as input. Tab. 4 and the supplement video results on real data show that event noise does not lead to a significant performance degradation, because NeRF training refers to multi-view input, which can inherently weaken the impact of sparse random noise in each individual view, as proved in RawNeRF [26]. Therefore, See-NeRF does not target the processing of event noise, which could be a future optimization direction of ERGB-based NVS methods.

### v2

@@ 2.2 ERGB-based novel view synthesis
Event cameras are bio-inspired sensors detecting brightness changes at individual pixels with high temporal resolution [12]. This scene radiance measurement paradigm has higher dynamic range and temporal resolution compared to traditional frame-based RGB cameras, attracting a lot of research in the fields of event-based image deblurring [33, 21, 26, 40, 47], HDR image reconstruction [42, 15, 44, 32, 45, 14, 8], and high-speed HDR video generation [37, 57, 48, 55]. Some recent works explore event-based NVS [20, 39, 28, 28, 29, 11, 53, 13, 51]. While other works integrate both image and event (ERGB) data into the deblurring NVS task, such as NeRF-based works [35, 34, 5, 36] and 3DGS-based works [10, 18, 50] Addressing the limitations of blurry LDR inputs, EvHDR-NeRF [6] first jointly exploits single-exposure images and events to reconstruct high-fidelity HDR radiance fields. However, these methods do not fully consider the gap between sensor output and physical radiance, leading to limited HDR and deblurring NeRF novel view synthesis performance.
@@ 5.1 Experiment settings
Synthetic dataset: The synthetic dataset consists of 8 HDR “Blender” scenes from HDR-NeRF [19]. Each scene contains 18 training views and 17 test views. We use Blender [1] and v2e [17] to generate blurry images with exposure time $\Delta t_{2}$ and corresponding events as in [35] for training views. We also provide sharp LDR images with exposure times $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ for the image-based HDR NVS methods without deblurring effects. For test views, we provide GT sharp HDR images and LDR images with $\Delta t_{1},\Delta t_{3}$ for HDR and novel exposure evaluation as in [19].
Real dataset: The real dataset includes 5 scenes captured under extreme lighting conditions using the DAVIS 346 event camera [41]. Each scene has 16 training views and 12 test views. For training views, we use a handheld DAVIS 346 to capture the blurry LDR images with exposure time $\Delta t_{2}$ along with the corresponding event data. We also captured sharp LDR images with a tripod with exposure times $\Delta t_{0},\Delta t_{2},\Delta t_{4}$. For the test views, we use a tripod to capture GT sharp LDR images with exposure times $\Delta t_{1}$, $\Delta t_{3}$. We also capture the GT sharp LDR images with $\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$, and use the classic multi-exposure HDR reconstruction algorithm DeBevec [9] to generate the GT sharp HDR image with the 5 sharp LDR images wtih different exposures for each test view as in RawHDR [56].
@@ 5.3 Deblurring NVS experiment
See-NeRFEvD eCRF [5] ours 22.02 .8669 .2474 16.43 .6932 .3210 29.52 .9267 .2776
@@ 5.4 Ablation study
Effect of event mapping field (EMF): Comparing See-NeRFNoEMF and See-NeRF in Tab. 3, our EMF marginally helps align the CRF curve (Fig. 8), achieving further improvement on HDR and Deblurring NVS. However, replacing our EMF with the eCRF model in EvDeblur-NeRF [5] leads to barely satisfactory reconstruction results and CRF estimation (See-NeRFEvD in Tab. 3 and Fig. 8).
@@ 5.5 Discussions
Robustness to LDR images with different exposures and event noise: As shown in Tab. 4, we conduct a robustness experiment on the synthetic “Catroom” scene for the HDR NVS task. Firstly, we use the single-exposure images with $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ as inputs, respectively. The results show that single-exposure inputs with longer exposure times yield better results, indicating that the low-light regions of the scene are more difficult to learn than the high-light regions. Secondly, we use events with different noise levels as input. Tab. 4 and the supplement video results on real data show that event noise does not lead to a significant performance degradation, because NeRF training refers to multi-view input, which can inherently weaken the impact of sparse random noise in each individual view, as proved in RawNeRF [30]. Therefore, See-NeRF does not target the processing of event noise, which could be a future optimization direction of ERGB-based NVS methods.
@@ D.3 Sensor-Physics grounded deblurring effect
Table 9: Discussion with EvHDR-GS
Imaging Model Event Real Data HDR Real Data Deblurring
Method Integration CRF Non-Idealities PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
EvHDR-GS N/A $\mathcal{F}(x)=x^{a}$ N/A 18.94 .7124 .2510 28.69 .9174 .2674
See-NeRF Eq. (6) MLP-Based RGB CRF Sec. 4.3 26.49 .8953 .1638 32.70 .9564 .1574
See-NeRF# N/A $\mathcal{F}(x)=x^{a}$ Sec. 4.3 19.75 .7239 .2696 28.35 .9242 .2657
D.4 Discussion with EvHDR-GS
A recent work, EvHDR-GS [7], uses 3D gaussian splatting as a backbone for the ERGB-based HDR video reconstruction task. Though it uses a similar design that RGB CRF is applied after volumetric rendering, it does not consider the image integration process and simply models the RGB CRF with one exponential function $\mathcal{F}(x)=x^{a}$ during the 2D mapping, while See-NeRF address this two aspects with Eq. (6) in the main paper and a three-independent MLP-Based RGB CRF model as shown in the Fig. 3 of the main paper. Besides, we model the event non-idealities with our event mapping fields. Therefore, EvHDR-GS underperforms See-NeRF on both deblurring and HDR tasks, and applying its imaging model to See-NeRF also degrades the performance significantly, as shown in Tab. 9.
D.5 Pose estimation
We use the event-enhanced pose estimation strategy of E2NeRF for our See-NeRF. Recent image-based deblurring NVS works provide different camera pose estimation strategies [43, 54, 24] and pose change estimation strategies for dynamic scenes [3, 25]. Integrating the event supervision into these advanced pose estimation strategies, as in EBAD-NeRF [36], would be a promising future research direction for the ERGB-based HDR deblurring NVS task.